OpenStudy (anonymous):

find the vertex -1/4x^2-1/3x

OpenStudy (amistre64):

--1/3 ----- = x vertex component 2(-1/4)

OpenStudy (amistre64):

1/3 ---- = -2/3 -1/2

OpenStudy (amistre64):

4/4(9) - -1(2)/(3)3 1/9 + 2/9 = 3/9 = 1/3 (-1/2, 1/3)